# ConsistencyDet: Robust Object Detector with Denoising Paradigm of Consistency Model

### v1

@@ 4 Experiments
In this section, we evaluate the performance of our model using two prevalent datasets: MS-COCO and LVIS, as cited in [27, 28]. Initially, we conduct experiments to ascertain the optimal values for key parameters within the MS-COCO and LVIS datasets, respectively. Subsequently, we compare the proposed ConsistencyDet framework against a range of established detection models, including the Diffusion Model. Finally, we undertake ablation studies to dissect and analyze the individual components of the ConsistencyDet framework.
[…]
4.1 Implementation details
@@
During the training phase, we initialize our model using pre-trained weights: those from ImageNet-1K for the ResNet backbone and ImageNet-21K [42] for the Swin-base backbone. The detection decoder, which is integrated into the DiffusionDet architecture, begins with Xavier initialization [43]. We employ the AdamW optimizer [44], with an initial learning rate set at $2.5\times{10}^{-5}$ and a weight decay parameter of ${10}^{-4}$. The training regime utilizes mini-batches of size 8, distributed across 4 GPUs. For MS-COCO, we adhere to a standard training schedule of 350,000 iterations, with the learning rate being reduced by a factor of ten at both 90,000 and 280,000 iterations. In the case of LVIS, similar reductions in the learning rate are scheduled at 100,000, 300,000, and 320,000 iterations. We apply augmentation techniques such as random horizontal flipping, scale jitter resizing (the shortest side ranges between 480 and 800 pixels and the longest side does not exceed 1333 pixels) and random crop augmentation. However, we abstain from using more robust augmentations like EMA, MixUp [45], or Mosaic [46].
4.2 Main Properties
The central characteristic of ConsistencyDet is its self-consistency, which ensures that the mapping effect from any point along the time axis back to the origin remains relatively stable. This stability signifies that once the model is adequately trained, it permits flexibility during inference by allowing adjustments to the number of sampling steps. By suitably increasing the sampling time steps, ConsistencyDet can improve its detection accuracy, though this enhancement comes at the expense of computational efficiency. Consequently, in diverse application contexts, the model offers the adaptability to adjust the number of sampling steps in response to specific requirements, thereby balancing between accuracy and algorithmic efficiency.
Dynamic boxes. The efficacy of ConsistencyDet is evaluated by comparing it with DiffusionDet and DETR on the MS-COCO dataset across a range of box quantities. As depicted in Fig. 5, the performance metrics for DiffusionDet and DETR are sourced from [21]. ConsistencyDet demonstrates a rapid improvement in performance as the number of boxes included in the evaluation escalates, with the gains stabilizing when $n_{p}>300$. The model attains optimal performance at $n_{p}=500$. In contrast, DETR exhibits peak AP at $n_{p}=300$, followed by a precipitous decline for $n_{p}>300$. Notably, at $n_{p}=4000$, DETR’s AP plummets to 26.4%, a 12.4% reduction from its peak AP of 38.8% achieved with 300 queries. Though DiffusionDet also experiences a marginal enhancement in performance with an increasing number of boxes, its results remain inferior to those of ConsistencyDet. Consequently, ConsistencyDet proves to be more robust across various scenarios with differing object counts, demonstrating superior noise resistance and transferability.
Figure 5: Performance comparisons with dynamic boxes. More proposal boxes in inference lead to accuracy improvement at first and tend towards stability for both ConsistencyDet and DiffusionDet, while cause a degradation in DETR. Furthermore, ConsistencyDet outperforms DiffusionDet for all these cases. Here, results of AP are all percentage data (%).
Progressive refinement. For both training and inference, the number of proposal boxes is fixed at 500 at each sampling time step. Within the MS-COCO dataset, ConsistencyDet demonstrates a stable performance enhancement as the number of sampling time steps increases from 1 to 10, as illustrated in Fig. 6. For instance, the AP for ConsistencyDet ascends from 45.73% at a single time step ($T\!\!=\!\!1$) to 46.88% at ten time steps ($T\!\!=\!\!10$), marking an absolute improvement of 1.15% in AP. This increment indicates that ConsistencyDet is capable of achieving higher accuracy with an optimal number of sampling time steps.
Figure 6: Performance analysis of progressive refinement. ConsistencyDet is evaluated with varied sampling time steps. It is trained on the MS-COCO dataset with ResNet-50 as the backbone, using 500 proposal boxes for training and evaluation at each sampling time step, and the number of total time steps $T=40$. Simultaneously, it is compared with DiffusionDet with same assignments, except $T=1000$.
Sampling step. As depicted in Fig. 7, the sampling phase starts with the initialization of noised boxes at the $t$-th time step, with the noise addition protocol. These noised boxes are then put into the model to generate predictions. The detection decoder predicts the category scores and box coordinates of current step. An initial filtration process ensues, whereby only the bounding boxes surpassing a predetermined confidence threshold are preserved. Under the theoretical guidance of Consistency Model, noised boxes are predicted, optimized and filtered for the next time step. At the end of each time step, additional proposals are supplemented by Box-renewal operation, ensuring the total amount of noise boxes reaches $n_{p}=500$ for the next iteration. In the final time step, NMS is applied to refine the results. By comparing the detection results in Fig. 7(f), it is evident that ConsistencyDet performs better than DiffusionDet, with more precise borders.
Figure 7: The comparison of visual reasoning process with one typical sampling step between ConsistencyDet and DiffusionDet. We selected a sample image from the COCO 2017 val set for testing with both ConsistencyDet and DiffusionDet. The initial noised boxes or verified boxes with low confidences are marked in write, while the boxes with high confidences are marked in red and the final predicted are marked in blue.
Parameter selection for significance. Given the disparity in dataset attributes between MS-COCO and LVIS [28], they exhibit distinct sensitivities to noise. For the optimization of ConsistencyDet on each dataset, it is essential to conduct comparative analyses to ascertain the ideal setting for the parameter $\sigma_{max}$. By calibrating this parameter, the model’s capacity to process noise can be fine-tuned, thereby facilitating its adaptation to the unique demands of each dataset and augmenting its efficacy across various contexts. This targeted adjustment plays a pivotal role in customizing the model’s performance for specific applications. The empirical outcomes, utilizing a ResNet-50 backbone with varying values of $\sigma_{max}$, are delineated in Table II.
TABLE II: Selecting the optimal $\sigma_{max}$ for COCO 2017 and LVIS v1.0 val set.
[…]
COCO LVIS
20 46.3 65.1 50.7 - - -
40 46.9 65.7 51.3 30.4 42.7 31.8
60 46.1 64.8 50.0 32.2 45.1 33.7
80 45.8 64.9 49.3 30.2 42.6 31.6
[…]
Bold fonts indicate the best performance.
4.3 Simulation analysis
The performance of ConsistencyDet is evaluated against other state-of-the-art detectors models [47, 13, 11, 10, 14] on the MS-COCO and LVIS datasets. The following provides an analysis of the simulation results.
TABLE III: Performance comparison of different object detectors on COCO 2017 val set.
[…]
@@ 4.3 Simulation analysis
RetinaNet 38.7 58.0 41.5 23.3 42.3 50.3
Faster R-CNN 40.2 61.0 43.8 24.2 43.5 52.0
Cascade R-CNN 44.3 62.2 48.0 26.6 47.7 57.7
DETR 42.0 62.4 44.2 20.5 45.8 61.1
Deformable DETR 43.8 62.6 47.7 26.4 47.1 58.0
Sparse R-CNN 45.0 63.4 48.2 26.9 47.2 59.5
DiffusionDet (1 step) 45.5 65.1 48.7 27.5 48.1 61.2
DiffusionDet (4 step) 46.1 66.0 49.2 28.6 48.5 61.3
DiffusionDet (8 step) 46.2 66.4 49.5 28.7 48.5 61.5
@@
RetinaNet 40.4 60.2 43.2 24.0 44.3 52.2
Faster R-CNN 42.0 62.5 45.9 25.2 45.6 54.6
Cascade R-CNN 45.5 63.7 49.9 27.6 49.2 59.1
DETR 43.5 63.8 46.4 21.9 48.0 61.8
Sparse R-CNN 46.4 64.6 49.5 28.3 48.3 61.6
DiffusionDet (1 step) 46.6 66.3 50.0 30.0 49.3 62.8
DiffusionDet (4 step) 46.9 66.8 50.4 30.6 49.5 62.6
DiffusionDet (8 step) 47.1 67.1 50.6 30.2 49.8 62.7
@@
Cascade R-CNN 51.9 70.9 56.5 35.4 55.2 51.9
Sparse R-CNN 52.0 72.2 57.0 35.8 55.1 52.0
DiffusionDet (1 step) 52.3 72.7 56.3 34.8 56.0 52.3
DiffusionDet (4 step) 52.7 73.5 56.8 36.1 56.0 68.9
DiffusionDet (8 step) 52.8 73.6 56.8 36.1 56.2 68.8
@@
Results of above evaluation metrics are all percentage data (%).
Bold fonts indicate the best performance while the italic fonts indicate the second best ones.
MS-COCO. The performance metrics of various detectors are compiled in Table III. ConsistencyDet, utilizing a ResNet-50 backbone, attains an AP of 46.9%, thereby outperforming established methodologies such as Faster R-CNN, RetinaNet, DETR, Sparse R-CNN, and DiffusionDet (with different sampling time steps). With the scaling up to a ResNet-101 backbone, ConsistencyDet’s performance is further enhanced, achieving an AP of 47.2%, which exceeds that of the aforementioned strong baselines. Moreover, when equipped with the Swin-base backbone [26] and pretrained on ImageNet-21k [42], ConsistencyDet reaches an impressive AP of 53.0%, eclipsing all the compared detectors. However, it should be noted that the proposed DiffusionDet exhibits the most robust overall performance, albeit with marginally lower results than other detectors on specific metrics.
TABLE IV: Performance comparison of different object detectors on LVIS v1.0 val set.
{threeparttable}
Method
AP
[…]
@@
Faster R-CNN
25.2
40.6
26.9
18.5
32.2
37.7
16.4
23.4
31.1
Cascade R-CNN
29.4
41.4
30.9
20.6
37.5
44.3
20.0
27.7
35.4
Sparse R-CNN
29.2
41.0
30.7
20.7
36.9
44.2
20.6
27.7
34.6
DiffusionDet(1step)
30.4
42.8
31.8
20.6
38.6
47.6
23.5
28.1
36.0
DiffusionDet(4step)
31.8
45.0
33.2
22.5
39.6
48.3 24.8
29.3
37.6
DiffusionDet(8step)
31.9 45.3
33.1
22.8 40.2 48.1
24.0
29.5 38.1
ConsistencyDet 32.2 45.1 33.7 23.1 40.4
47.7
24.3 30.1 38.0
@@
Faster R-CNN
27.2
42.9
29.1
20.3
35
40.4
18.8
25.4
31.1
Cascade R-CNN
31.6
43.8
33.4
22.3
39.7
47.3
23.9
29.8
35.4
Sparse R-CNN
30.1
42.0
31.9
21.3
38.5
45.6
23.5
27.5
34.6
DiffusionDet(1step)
31.9
44.6
33.1
21.6
40.3
49.0
23.4
30.5
36.0
DiffusionDet(4step)
32.9
46.5
34.3
23.3
41.1
49.9
24.2
31.3
37.6
DiffusionDet(8step)
33.5 47.3 34.7 23.6 41.9 49.8 24.8 32.0 39.0
ConsistencyDet 33.1
46.1
34.8
22.9
41.5
49.4
24.9
31.2
38.7
@@
DiffusionDet(1step)
40.6
54.8
42.7
28.3
50.0
61.6
33.6
39.8
44.6
DiffusionDet(4step)
41.9
57.1
44.0
30.3
50.6
62.3
34.9
40.7
46.3
DiffusionDet(8step)
42.1 57.8 44.3 31.0 51.3 62.5
34.3
41.0 46.7
ConsistencyDet 42.4 57.1 44.9 30.3 51.5 63.4 34.9 41.5 46.8
{tablenotes}
Results of above evaluation metrics are all percentage data (%).
@@
Bold fonts indicate the best performance while the italic fonts indicate the second best ones.
LVIS v1.0. Table IV presents a comparative analysis of object detection performance between ConsistencyDet and various other detection models. Reproductions of Faster R-CNN and Cascade R-CNN have yielded AP scores of 25.2% and 27.2% with a ResNet-50 backbone, and 29.4% and 31.6% with a ResNet-101 backbone, respectively. ConsistencyDet, employing a ResNet-50 backbone, has achieved an AP of 32.2%, thereby exceeding the performance of established methods such as Faster R-CNN, RetinaNet, DETR, and Sparse R-CNN, while also remaining competitive with DiffusionDet. As the backbone is scaled up, ConsistencyDet exhibits consistent performance enhancements. With the ResNet-101 backbone, ConsistencyDet reaches an AP of 33.1%, surpassing robust baselines including Faster R-CNN, RetinaNet, DETR, Sparse R-CNN, and DiffusionDet at a quarter step, though it falls slightly behind DiffusionDet at an eight-step iteration. Furthermore, when leveraging the Swin-base backbone [26] pretrained on ImageNet-21k [42], ConsistencyDet achieves a significant AP of 42.4%, outstripping all the aforementioned baselines and establishing new benchmarks for models such as Faster R-CNN, RetinaNet, DETR, Sparse R-CNN, and DiffusionDet.
4.4 Ablation Study
@@
Comprehensive ablation studies were conducted to elucidate the characteristics of the proposed ConsistencyDet model on the MS-COCO and LVIS v1.0 datasets. These simulations utilized a ResNet-50 architecture equipped with a FPN as the primary backbone, with no further modifications or enhancements specified.
[…]
It is important to note that the optimal threshold selection may exhibit slight variations when coupled with different backbone architectures within the dataset, as evidenced in the middle column of Table VI. To ensure the robustness and generalizability of the code during the testing phase, unified parameters were established, and the corresponding AP is documented in the right column of Table VI.
TABLE V: Optimal choices for the thresholds for the Box-renewal and NMS operation.
[…]
@@
[…]
TABLE VI: Optimal choices for Box-renewal threshold and score threshold across different datasets and backbones.
{threeparttable}
[…]
COCO ResNet-50 0.98 0.64 46.9 0.98 0.64 46.9
ResNet-101 0.85 0.60 47.2 0.98 0.64 47.0
Swin-Base 0.98 0.62 53.0 0.98 0.64 52.8
LVIS ResNet-50 0.85 0.60 32.2 0.85 0.60 32.2
ResNet-101 0.85 0.60 33.1 0.85 0.60 33.1
Swin-Base 0.85 0.60 42.4 0.85 0.60 42.4
{tablenotes}
@@
Bold fonts indicate the best performance.
Figure 9: Typical instances of detected objects by the proposed ConsistencyDet. Their objects in these images are detected with correct categories and accurate borders.
Training strategy. In Algorithm 4, we mentioned two different training strategies, distinguished by whether to utilize a pre-trained DiffusionDet model for distillation training of ConsistencyDet. In Fig. 8, the loss curves of these two strategies are compared in the coordinate system. Both of them are trained with 4 GPUs, and the batch size is set to be 2 for each GPU. The training process adopts Res-Net50 as the backbone and is conducted on MS-COCO dataset. While DiffusionDet could, in theory, achieve better results with an adequate number of denoising time steps, the complexity of its operational steps presents challenges in effective training, which in turn limits its performance in practice. Hence, our experimental efforts primarily focused on the training strategy of the Consistency Model in isolation.
From Fig. 8 (b) and (c), it is evident that employing distillation training can effectively mitigate the issue of outliers encountered during the training process. At the latter half of the timeline, the stochastic nature introduced by noise may lead to unreasonable values of box coordinates sometimes, where the top-left coordinates are smaller than the bottom-right coordinates. This illogical issue results in negative GIoU loss values, potentially causing the model to converge in an erroneous direction, which could pose potential risks. Therefore, the introduction of distillation brings corresponding benefits.
Since the predictions of current DiffusionDet still have a big gap with GT information, if the position information of the boxes predicted by DiffusionDet is used as the reference information from the origin of the timeline, this operation inevitably leads to inadequate guidance in the training process of ConsistencyDet, and results in a larger total loss with poor performance, shown as Fig. 8 (a). Therefore, considering comprehensively, the training process of this work adopts the independent training strategy, which has achieved better results. Also, the aforementioned illogical issue can be restrained with the limitation of value ranges conservatively.
Accuracy vs. speed. Table VII compares the inference speeds of ConsistencyDet and DiffusionDet, both employing a ResNet-50 backbone, on the COCO 2017 validation set. The runtime performance was measured using a single NVIDIA RTX 3080 GPU and a batch size of one. DiffusionDet’s evaluation included various sampling time steps ($t_{ss}=20,40,60,100$) with a total time step of $T=1000$, from which AP and Frames Per Second (FPS) were recorded. ConsistencyDet was tested at total time steps of $t_{ss}=2$ and $10$ with $T=40$. The experimental outcomes reveal that ConsistencyDet not only achieves a substantial enhancement in inference speed when compared to DiffusionDet but also demonstrates a modest improvement of 0.4% in AP.
TABLE VII: Accuracy vs. speed comparison of ConsistencyDet and DiffusionDet at different sampling time steps.
{threeparttable}
Model $T$ $n_{ss}$ AP FPS
DiffusionDet 1000 20 46.5 0.954
1000 40 46.4 0.473
1000 60 46.5 0.309
ConsistencyDet 40 2 46.8 6.892
40 10 46.9 1.783
{tablenotes}
Results of AP are all percentage data (%).
Bold fonts indicate the best performance while the italic fonts indicate the second best ones.
In this study, we introduce a novel object detection framework that incorporates the Consistency Model, which is compatible with commonly utilized backbone and head architectures. This framework does not impose special requirements on the network structures. The proposed ConsistencyDet has demonstrated superior overall performance compared to other well-established detectors. Fig. 9 showcases visualizations of typical instances where ConsistencyDet has accurately detected objects with precise boundaries and correct category annotations. Although there is still a performance gap when compared to top-performing detectors cited in references [48, 49, 50], we anticipate that the performance of DiffusionDet could be further enhanced through its integration with specially designed architectures.
Despite its promising performance, ConsistencyDet is not without limitations. Its detection accuracy for larger objects is marginally lower than that of DiffusionDet and DETR when using certain backbones, partly due to the noise addition strategy not fully taking into account the scale of the objects. While it operates faster than DiffusionDet, there remains a need for further efficiency improvements to meet the demands of real-time applications.
5 Conclusions

### v5

@@ IV Experiment
In this section, we evaluate the performance of our model using two prevalent datasets: MS-COCO and LVIS, as cited in [24, 25]. Initially, we conduct experiments to ascertain the optimal values for key parameters in the MS-COCO and LVIS datasets. Subsequently, we compare the proposed ConsistencyDet framework against a range of established detection models, including the Diffusion Model. Finally, we undertake ablation studies to dissect and analyze the individual components of the ConsistencyDet framework.
MS-COCO [24] comprises 118,000 training images in the train2017 subset and 5,000 validation images in the val2017 subset, spanning 80 different object categories. A performance evaluation is conducted using the standard metrics of box average precision (AP), and AP at IoU thresholds of 0.5 (AP50) and 0.75 (AP75). The IoU metric quantifies the proportionate overlap between each predicted bounding box and the corresponding GT boxes, offering important insight into the precision of the object detector at various levels of accuracy. Besides, scale-specific AP scores are also adopted: APs / APm / APl for small / median / large objects separately.
LVIS v1.0 [25] encompasses 100,000 training images and 20,000 validation images, drawing from the same source images as MS-COCO. It focuses on large-vocabulary object detection and instance segmentation by annotating a long-tailed distribution of objects across 1,203 categories. The evaluation framework for this dataset incorporates a suite of metrics: AP / AP50 / AP75, and APs / APm / APl for small / median / large objects separately (scale-specific AP scores), APr / APc / APf for rare / common / frequent categories separately (category-specific AP scores). These metrics provide a multifaceted evaluation of model performance across varying levels of detection precision and category frequency.
IV-A Implementation detail
@@
During training, we initialize our model using pre-trained weights from ImageNet-1K for the ResNet backbone and ImageNet-21K [38] for the Swin-base backbone. The detection decoder in the ConsistencyDet architecture starts with Xavier initialization [39]. We utilize the AdamW optimizer [40], with an initial learning rate of $2.5\times 10^{-5}$ and a weight decay of $10^{-4}$. Training employs mini-batches of size 8 across 4 GPUs. For MS-COCO, we follow a standard training schedule of 350,000 iterations, reducing the learning rate by a factor of ten at 90,000 and 280,000 iterations. For LVIS, reductions occur at 100,000, 300,000, and 320,000 iterations. We apply augmentation techniques, including random horizontal flipping, scale jitter resizing (with the shortest side between 480 and 800 pixels and the longest side not exceeding 1333 pixels), and random cropping. However, we do not utilize more robust augmentations like EMA, MixUp [41], or Mosaic [42].
TABLE II: Performance comparison of different object detectors on the COCO 2017 val set.
Method AP AP50 AP75 APs APm APl
@@
Faster R-CNN [43] 40.2 61.0 43.8 24.2 43.5 52.0
Cascade R-CNN [44] 44.3 62.2 48.0 26.6 47.7 57.7
DETR [13] 42.0 62.4 44.2 20.5 45.8 61.1
Sparse R-CNN [14] 45.0 63.4 48.2 26.9 47.2 59.5
DiffusionDet [19] 46.2 66.4 49.5 28.7 48.5 61.5
@@
Faster R-CNN [43] 42.0 62.5 45.9 25.2 45.6 54.6
Cascade R-CNN [44] 45.5 63.7 49.9 27.6 49.2 59.1
DETR [13] 43.5 63.8 46.4 21.9 48.0 61.8
Sparse R-CNN [14] 46.4 64.6 49.5 28.3 48.3 61.6
DiffusionDet [19] 47.1 67.1 50.6 30.2 49.8 62.7
@@
Cascade R-CNN [44] 51.9 70.9 56.5 35.4 55.2 51.9
Sparse R-CNN [14] 52.0 72.2 57.0 35.8 55.1 52.0
DiffusionDet [19] 52.8 73.6 56.8 36.1 56.2 68.8
@@
  • 1
    Results of the above evaluation metrics are all percentage data (%).
  • 2
    Bold font indicates the best performance while underlined font indicates the second best.
TABLE III: Performance comparison of different object detectors on LVIS v1.0 val set.
Method AP AP50 AP75 APs APm APl
@@
Faster R-CNN [43] 25.2 40.6 26.9 18.5 32.2 37.7
Cascade R-CNN [44] 29.4 41.4 30.9 20.6 37.5 44.3
Sparse R-CNN [14] 29.2 41.0 30.7 20.7 36.9 44.2
DiffusionDet [19] 31.9 45.3 33.1 22.8 40.2 48.1
ConsistencyDet 32.2 45.1 33.7 23.1 40.4 47.7
@@
Faster R-CNN [43] 27.2 42.9 29.1 20.3 35 40.4
Cascade R-CNN [44] 31.6 43.8 33.4 22.3 39.7 47.3
Sparse R-CNN [14] 30.1 42.0 31.9 21.3 38.5 45.6
DiffusionDet [19] 33.5 47.3 34.7 23.6 41.9 49.8
ConsistencyDet 33.1 46.1 34.8 22.9 41.5 49.4
@@
DiffusionDet [19] 42.1 57.8 44.3 31.0 51.3 62.5
ConsistencyDet 42.4 57.1 44.9 30.3 51.5 63.4
  • 1
    Results of the above evaluation metrics are all percentage data (%).
  • 2
@@
    Bold font indicates the best performance while underlined font indicates the second best performance.
IV-B Main Property
The key characteristic of ConsistencyDet is self-consistency, which stabilizes the mapping from any point along the time axis back to the origin. This stability provides flexibility during inference, allowing for adjustments in the number of sampling steps. Increasing the sampling steps can improve detection accuracy, although it may reduce computational efficiency. Consequently, the model can balance accuracy and efficiency based on specific application needs. Additionally, ConsistencyDet exhibits strong noise resistance, allowing for dynamic adjustments in the number of bounding boxes during inference. Some detection results are shown in Fig. 8.
Refinement Through Iterative Sampling. For inference, the number of proposal boxes is fixed at 500 at each sampling step. In the MS-COCO dataset, ConsistencyDet shows stable performance improvements as sampling steps increase from 1 to 10, as illustrated in Fig. 5. The AP for ConsistencyDet rises from 45.73% at $T=1$ to 46.88% at $T=10$, indicating that higher accuracy can be achieved with optimal sampling steps.
Figure 5: Performance analysis of progressive refinement. ConsistencyDet is evaluated with varied sampling time steps. It is trained on the MS-COCO dataset with ResNet-50 as the backbone, using 500 proposal boxes for evaluation at each sampling time step. Simultaneously, the accuracy of ConsistencyDet surpasses that of DiffusionDet under the same experimental parameters. Here, the results of AP are all percentage data (%).
Dynamic Box Sampling for Improved Robustness. The efficacy of ConsistencyDet is evaluated against DiffusionDet and DETR on the MS-COCO dataset, as shown in Fig. 6. ConsistencyDet improves performance significantly as the number of boxes increases, stabilizing when $n_{p}>300$ and achieving optimal performance at $n_{p}=500$. In contrast, DETR peaks at $n_{p}=300$, then declines sharply, dropping to 26.4% at $n_{p}=4000$, a 12.4% reduction from its peak. At the same time, ConsistencyDet outperforms DiffusionDet for all cases. Moreover, this experimental results demonstrate that ConsistencyDet significantly surpasses traditional detection methods like DETR in terms of noise resistance.
Figure 6: Performance comparison of dynamic box sampling for improved robustness. More proposal boxes in inference lead to accuracy improvement at first and tend towards stability for both ConsistencyDet and DiffusionDet, while causing a degradation in DETR due to its lack of noise resistance. Furthermore, ConsistencyDet outperforms DiffusionDet for all these cases. Here, the results of AP are all percentage data (%).
Sampling step. As depicted in Fig. 7, the sampling phase starts with the initialization of noised boxes at the $t$-th time step, with the noise addition protocol. These noised boxes are then put into the model to generate predictions. The detection decoder predicts the category scores and box coordinates of the current step. An initial filtration process ensues, whereby only the bounding boxes surpassing a predetermined confidence threshold are preserved. Under the theoretical guidance of the Consistency Model, noised boxes are predicted, optimized and filtered for the next time step. At the end of each time step, additional proposals are supplemented by the Box-renewal operation, ensuring the total amount of noise boxes reaches $n_{p}=500$ for the next iteration. In the final time step, NMS is applied to refine the results. By comparing the detection results in Fig. 7(f), it is evident that ConsistencyDet performs better than DiffusionDet, with more precise borders.
Figure 7: A comparison of the visual reasoning process with one typical sampling step between ConsistencyDet and DiffusionDet. We selected a sample image from the COCO 2017 val set for testing with both ConsistencyDet and DiffusionDet. The initial noised boxes or verified boxes with low confidence are marked in white, while the boxes with high confidence are marked in red and the final predictions are marked in blue.
IV-C Simulation analysis
The performance of ConsistencyDet is evaluated against other state-of-the-art detector models [43, 44, 14, 13, 19] on the MS-COCO and LVIS v1.0 datasets. The following provides an analysis of the simulation results.
MS-COCO. Performance metrics for different detectors are summarized in table II. ConsistencyDet, with a ResNet-50 backbone, achieves an AP of 46.9%, outperforming Faster R-CNN, DETR, Sparse R-CNN, and DiffusionDet. Upgrading to a ResNet-101 backbone boosts ConsistencyDet’s AP to 47.2%, exceeding these baselines. Additionally, with a Swin-base backbone [26] pretrained on ImageNet-21k [38], ConsistencyDet achieves an impressive AP of 53.0%, outperforming all compared detectors. This shows our method’s consistent performance improvement with varied backbones, confirming its effectiveness in detection accuracy.
Figure 8: Typical instances of detected objects by the proposed ConsistencyDet. The objects in these images are detected with correct categories and accurate borders.
LVIS v1.0.  table III compares the object detection performance of ConsistencyDet with several other models. Using a ResNet-50 backbone, reproductions of Faster R-CNN and Cascade R-CNN yielded APs of 25.2% and 29.4%, respectively, while ConsistencyDet achieved 32.2%, outperforming Faster R-CNN, DETR, and Sparse R-CNN and remaining highly competitive with DiffusionDet. With the ResNet-101 backbone, ConsistencyDet reaches an AP of 33.1%, slightly below DiffusionDet but above other baselines. When using the Swin-base backbone [26], ConsistencyDet achieves 42.4% AP, outperforming all other mentioned baselines.
IV-D Ablation Study
@@
Comprehensive ablation studies were carefully conducted to further elucidate the key characteristics of the proposed ConsistencyDet model on the MS-COCO and LVIS v1.0 datasets. These simulations utilized a ResNet-50 architecture equipped with a FPN as the primary backbone, with no additional modifications or enhancements specified.
Box Renewal threshold and NMS threshold. The left column of table IV illustrates the impact of the score threshold, denoted as $B_{th}$, on the Box-renewal process. A threshold value of 0.0 implies that Box-renewal is not employed. According to the evaluation on the COCO 2017 validation set, employing a threshold of 0.98 marginally outperforms other threshold settings.The right column of table IV delineates the effects of varying the NMS score threshold, represented as $N_{th}$, on the AP. Setting the threshold to 0.0 signifies the absence of any NMS procedure. An analysis of the COCO 2017 validation set suggests that a threshold of 0.64 yields a performance that is modestly superior relative to the other thresholds investigated.
TABLE IV: Optimal choices for the thresholds for the Box-renewal and NMS operation.
$B_{th}$ AP AP50 AP75 $N_{th}$ AP AP50 AP75
@@
Training strategy. Algorithm 4 provides two training strategies, differentiated by whether to utilize a pre-trained DiffusionDet model for the distillation training of ConsistencyDet. In Fig. 9, the loss curves of these two strategies are compared in the coordinate system. Both are trained with 4 GPUs, and the batch size is set to 2 for each GPU. The training process adopts Res-Net50 as the backbone and is conducted on the MS-COCO dataset. While DiffusionDet could, in theory, achieve better results with an adequate number of denoising time steps, the complexity of its operational steps presents challenges in effective training, which in turn limits its performance in practice. Hence, our experimental efforts primarily focus on the training strategy of the Consistency Model in isolation. From Fig. 9(b)-(c), it is evident that employing distillation training can effectively mitigate the issue of outliers encountered during the training process. At the latter half of the timeline, the stochastic nature introduced by noise may sometimes lead to unreasonable values of box coordinates, where the top-left coordinates are smaller than the bottom-right coordinates. This illogical issue results in negative GIoU loss values, potentially causing the model to converge in an erroneous direction, which could pose potential risks. Therefore, the introduction of distillation brings corresponding benefits. Since the predictions of current DiffusionDet still have a big gap with GT information, if the position information of the boxes predicted by DiffusionDet is used as the reference information from the origin of the timeline, this operation inevitably leads to inadequate guidance in the training process of ConsistencyDet, and results in a larger total loss with poor performance, shown as Fig. 9(a). Therefore, after comprehensive consideration, this work adopts the independent training strategy for the training process which achieves better results. Also, the aforementioned illogical issue can be restrained with the conservative limitation of value ranges.
Accuracy vs Speed. The inference speeds of ConsistencyDet and DiffusionDet, both utilizing a ResNet-50 backbone, are compared in table V. Runtime performance was measured on a single RTX3080 GPU with a batch size of one. DiffusionDet’s evaluation includes sampling timesteps ($t_{ss}\!\!=\!\!2/10/20/40$), recording both AP and FPS. ConsistencyDet was tested at total timesteps $t_{ss}=2/10/20$. The experimental results show that ConsistencyDet reaches its optimal detection accuracy at $t_{ss}=10$, while at $t_{ss}=2$, its accuracy is only 0.1% lower than the peak performance, aligning well with the “few-step denoising” mechanism of ConsistencyDet. In contrast, DiffusionDet requires $t_{ss}=20$ to achieve its best detection accuracy, but even at this setting, it performs worse than ConsistencyDet at $t_{ss}=2$ and is 7.2 times less efficient in inference speed. This difference clearly demonstrates the significant advantages of ConsistencyDet in both speed and accuracy. Moreover, ConsistencyDet consistently maintains superior detection accuracy at all sampling steps compared to DiffusionDet, further underscoring its stability. In summary, our method achieves stable and enhanced detection performance with only a few sampling steps, overcoming the main drawback of DiffusionDet, which sacrifices inference speed to improve detection accuracy.
TABLE V: Accuracy vs. speed comparison of ConsistencyDet and DiffusionDet at different sampling time steps.
Model $n_{ss}$ AP FPS
DiffusionDet 2 45.82 6.734
10 46.33 1.987
20 46.54 0.954
40 46.44 0.473
ConsistencyDet 2 46.8 6.892
10 46.88 1.783
20 46.87 0.845
  • 1
@@
  • 2
    Bold font indicates the best performance.
V Conclusion